Find the amount if $500 is invested at 15% compounded monthly for 17 months.
Amount = $

Answer by stanbon(75887)

Find the amount if $500 is invested at 15% compounded monthly for 17 months.
---------
A(t) = P(1+(r/n))^(nt)
---
A(17/12) = 500(1+(0.15/12))^(12(17/12)
-----
A(17/12) = 500(1.0125)^(17)= $617.57
------------------------
Cheers,
Stan H.
=================